# A Malliavin-Gamma calculus approach to Score Based Diffusion Generative models for random fields

### v1

@@ 3. Denoising via abstract time-reversal
$\mathbb{E}_{\mathfrak{m}}\biggl{[}\frac{\Gamma(\rho_{t},\,u)}{\rho_{t}}(X)\,v(X)\biggr{]}=\mathbb{E}_{\mathfrak{m}}\biggl{[}\langle\bar{\nabla}\rho_{t}(X),\,h\rangle_{\mathcal{H}_{\mathfrak{m}}}\,\frac{v(X)}{\rho_{t}(X)}\biggr{]}=e^{-\frac{t+\varepsilon}{2}}\,\mathbb{E}_{\mathfrak{m}}\biggl{[}\langle P_{t-\varepsilon}\bar{\nabla}\rho_{\varepsilon}(X),\,h\rangle_{\mathcal{H}_{\mathfrak{m}}}\,\frac{v(X)}{\rho_{t}(X)}\biggr{]}$
@@
where the last step follows from the fact that $P_{t-\varepsilon}$ is self-adjoint in $\mathrm{L}^{2}(\mathfrak{m})$. Next, if $\delta$ denotes the divergence operator associated to $\bar{\nabla}$ (i.e., the $\mathrm{L}^{2}(\mathfrak{m})$-adjoint of the operator $\bar{\nabla}$ [29, Section 1.3]), then from [29, Lemma 1.3.1 and Proposition 1.3.3] we have
@@
(3.4) $\lim_{\varepsilon\downarrow 0}\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl{[}I(h)(X_{\varepsilon})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr{]}=\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl{[}I(h)(X_{0})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr{]}=\mathbb{E}_{X\sim\overrightarrow{\mathrm{P}}_{t}}\biggl{[}\frac{v(X)}{\rho_{t}(X)}\,\mathbb{E}_{\overrightarrow{\mathrm{P}}}[I(h)(X_{0})|X_{t}=X]\biggr{]}$
@@
Let us now focus on the last term appearing in (3.3). By relying again on the commutation estimate Section 3, the self-adjointness of the semigroup and on the properties of the divergence operator $\delta$ (cf. [29, Lemma 1.3.1 and Proposition 1.3.3]), as already done before, we have
@@
$\mathbb{E}_{\mathfrak{m}}\biggl{[}\frac{\Gamma(\rho_{t},\,u)}{\rho_{t}}(X)\,v(X)\biggr{]}=e^{-\nicefrac{{t}}{{2}}}\,\mathbb{E}_{\mathfrak{m}}[v(X)\,\mathbb{E}_{\overrightarrow{\mathrm{P}}}[I(h)(X_{0})|X_{t}=X]]-e^{-t}\mathbb{E}_{\mathfrak{m}}[v(X)\,I(h)(X)]$
$+e^{-t}\,\mathbb{E}_{\mathfrak{m}}\biggl{[}\frac{\Gamma(\rho_{t},\,u)}{\rho_{t}}(X)\,v(X)\biggr{]}\,,$
@@
The previous proof highlights the importance of having a commutation relation between the semigroup and the Malliavin gradient. Indeed, this is the only step in the proof where we have explicitly relied on the OU nature of the forward dynamics. This suggests that the approach presented here might be generalized to different noising (diffusion) dynamics, provided a commutation relation of the form ¯$\bar{\nabla}P_{t}=f(t)P_{t}\bar{\nabla}$ for a general $f(t)\in(0,1)$. Particularly, this further highlights the fact that the score equals a conditional expectation of $X_{0}$ (conditioned to $\overrightarrow{X}_{t}$) provided the commutation between Malliavin gradient and semigroup, and henceforth its validity might be extended beyond the OU dynamics.
@@
For any $u\in\mathrm{L}^{1}(\mathrm{H},\mathfrak{m})$ and any $t\geq 0$ we have
¯$\bar{\nabla}P_{t}\,u=e^{-\nicefrac{{t}}{{2}}}P_{t}\bar{\nabla}u\,.$
Proof.
If $u\in\mathcal{PC}$ the thesis trivially follows from the explicit expressions (2.5) combined with [29, Lemma 1.4.2]. The extension to $\mathrm{L}^{1}(\mathrm{H},\mu)$ follows from the density of $\mathcal{PC}\subseteq\mathrm{L}^{1}(\mathrm{H},\mathfrak{m})$ and the closability of $\bar{\nabla}$ [28], combined with the fact that $(P_{t})_{t\geq 0}$ is a linear contraction operator on $\mathrm{L}^{1}(\mathrm{H},\mu)$ [28, Proposition 2.8.6]. ∎
[…]
(3.6) $\begin{cases}\mathrm{d}\overleftarrow{X}_{s}=-\frac{1}{2}\,\overleftarrow{X}_{s}\mathrm{d}s+\sigma_{T-s}^{-1}\,\mathbb{E}_{\overrightarrow{\mathrm{P}}}[\overrightarrow{X}_{0}-e^{\nicefrac{{T-s}}{{2}}}\overrightarrow{X}_{T-s}|\overrightarrow{X}_{T-s}=\overleftarrow{X}_{s}]\,\mathrm{d}s+\mathrm{d}W^{\mathcal{H}_{\mathfrak{m}}}_{s}\\ \overleftarrow{X}_{0}\sim\overrightarrow{\mathrm{P}}_{T}\,,\end{cases}$
@@
Claim: $\mathrm{Q}^{J}=\overleftarrow{\mathrm{P}}^{J}$. Let us start by showing that $\mathrm{Q}^{J}$ is a solution to the (finite-dimensional) martingale problem associated to the generator
(3.7) $\overleftarrow{L}_{s}^{J}u(x)=\overrightarrow{L}^{J}u(x)+\sigma_{T-s}^{-1}\sum_{j=1}^{N}\partial_{\ell_{j}}u(x)\,\mathbb{E}_{\overrightarrow{\mathrm{P}}^{J}}[X_{0}^{\ell_{j}}-e^{-\nicefrac{{T-s}}{{2}}}X_{T-s}^{\ell_{j}}|X_{T-s}^{J}=x],\quad\forall\,x\in\mathbb{R}^{N},\,\forall\,u\in\mathcal{C}^{\infty}_{\mathrm{b}}(\mathbb{R}^{N}),$
@@
$\int_{0}^{t}\sigma_{T-s}^{-1}\,\mathbb{E}_{\mathrm{Q}}\biggl{[}\mathbb{E}_{\overrightarrow{\mathrm{P}}}[I(\bar{\nabla}u^{\mathrm{ext}}(\overrightarrow{X}_{s}))(\overrightarrow{X}_{0}-e^{-\nicefrac{{T-s}}{{2}}}\overrightarrow{X}_{T-s})|\overrightarrow{X}_{T-s}=X_{s}]\biggr{]}\mathrm{d}s$
=(2.4)${=}\int_{0}^{t}\mathbb{E}_{\mathrm{Q}}\biggl{[}-\frac{1}{2}\nabla u(X^{J}_{s})\cdot X^{J}_{s}+\frac{1}{2}\sum_{j=1}^{N}C_{\ell_{j}}\partial^{2}_{jj}u(X_{s}^{J})\biggr{]}\mathrm{d}s$
$+\int_{0}^{t}\sigma_{T-s}^{-1}\mathbb{E}_{\mathrm{Q}}\biggl{[}\nabla u(X^{J}_{s})\cdot\mathbb{E}_{\overrightarrow{\mathrm{P}}}[\overrightarrow{X}^{J}_{0}-e^{\nicefrac{{T-s}}{{2}}}\overrightarrow{X}_{T-s}^{J}|\overrightarrow{X}_{T-s}=X_{s}]\biggr{]}\mathrm{d}s$
$=\int_{0}^{t}\mathbb{E}_{\mathrm{Q}}\overrightarrow{L}^{J}u(X^{J}_{s})\mathrm{d}s+\int_{0}^{t}\sigma_{T-s}^{-1}\mathbb{E}_{\mathrm{Q}}\biggl{[}\nabla u(X^{J}_{s})\cdot\mathbb{E}_{\overrightarrow{\mathrm{P}}^{J}}[\overrightarrow{X}^{J}_{0}-e^{\nicefrac{{T-s}}{{2}}}\overrightarrow{X}_{T-s}^{J}|\overrightarrow{X}_{T-s}^{J}=X_{s}^{J}]\biggr{]}\mathrm{d}s\,,$
@@
4. $\Gamma$-Score approximation, sampling scheme and convergence results
@@
Therefore the core of any generative model (based on our OU noising forward dynamics) boils down to learning from nosing forward dynamics the conditional expectation function $\mathbb{E}_{\overrightarrow{\mathrm{P}}}[\overrightarrow{X}_{0}|\overrightarrow{X}_{T-s}=\Phi]$. Owing to the characterization of conditional expectations as $\mathrm{L}^{2}$-projections [19, Corollary 8.17], we know that for any $\sigma(\overrightarrow{X}_{T-s})$-measurable random field $Y\in\mathrm{L}^{2}(\overrightarrow{\mathrm{P}}_{T-s})$
@@
(4.1) $\min_{\theta\in\Theta}\,\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\,\|\overrightarrow{X}^{(i)}_{0}-\mathrm{v}_{T-t_{j}}^{\theta}(\overrightarrow{X}^{(i)}_{T-t_{j}})\|_{\mathrm{H}}^{2}\,,$
@@
Input: $N$ samples $X^{(i)}\sim\mu_{\mathrm{data}}$ from our data distribution, time horizon $T\gg 1$ large enough, partition $0=t_{0}<t_{1}<\cdots<t_{M}=T$ (usually uniformly at random or equidistributed), parametrized vector field family $\{\mathrm{v}^{\theta}:\theta\in\Theta\}$
@@ 2       Initialize $\overrightarrow{X}_{0}^{(i)}=X^{(i)}$
(4.2) $\begin{cases}\mathrm{d}Y_{s}=-\frac{1}{2}\,Y_{s}\mathrm{d}s+\sigma_{T-t_{j}}^{-1}\,(\mathrm{v}^{\theta^{\star}}_{T-t_{j}}(Y_{t_{j}})-e^{\nicefrac{{T-t_{j}}}{{2}}}Y_{t_{j}})\mathrm{d}s+\mathrm{d}W^{\mathcal{H}_{\mathfrak{m}}}_{s}\quad\forall\,s\in[t_{j},t_{j+1}]\\ Y_{0}\sim\mathfrak{m}\,,\end{cases}$

### v2

@@ 3. Denoising via abstract time-reversal
$\mathbb{E}_{\mathfrak{m}}\biggl[\frac{\Gamma(\rho_{t},\,u)}{\rho_{t}}(X)\,v(X)\biggr]=$ $\,\mathbb{E}_{\mathfrak{m}}\biggl[\langle\bar{\nabla}\rho_{t}(X),\,h\rangle_{\mathcal{H}_{\mathfrak{m}}}\,\frac{v(X)}{\rho_{t}(X)}\biggr]$
$=$ $\,e^{-\frac{t+\varepsilon}{2}}\,\mathbb{E}_{\mathfrak{m}}\biggl[\langle P_{t-\varepsilon}\bar{\nabla}\rho_{\varepsilon}(X),\,h\rangle_{\mathcal{H}_{\mathfrak{m}}}\,\frac{v(X)}{\rho_{t}(X)}\biggr]$
@@
where the last step follows from the fact that $P_{t-\varepsilon}$ is self-adjoint in $\mathrm{L}^{2}(\mathfrak{m})$. Next, if $\delta$ denotes the divergence operator associated to $\bar{\nabla}$ (i.e., the $\mathrm{L}^{2}(\mathfrak{m})$-adjoint of the operator $\bar{\nabla}$ [34, Section 1.3]), then from [34, Lemma 1.3.1 and Proposition 1.3.3] we have
@@
(3.4) $\lim_{\varepsilon\downarrow 0}\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl[I(h)(X_{\varepsilon})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr]$ $\,=\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl[I(h)(X_{0})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr]$
$=$ $\,\mathbb{E}_{X\sim\overrightarrow{\mathrm{P}}_{t}}\biggl[\frac{v(X)}{\rho_{t}(X)}\,\mathbb{E}_{\overrightarrow{\mathrm{P}}}[I(h)(X_{0})|X_{t}=X]\biggr]$
@@
To be more precise, in order to consider the vanishing $\varepsilon$ limit above, we proceed as follows. Firstly, we rely on the properties of the divergence operator $\delta$ (cf. [34, Lemma 1.3.1 and Proposition 1.3.3]) and write
$\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl[I(h)(X_{\varepsilon})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr]=$ $\,\mathbb{E}_{X\sim\mathfrak{m}}\biggl[\rho_{\varepsilon}(X)\,I(h)(X)\,\mathbb{E}_{\mathrm{R}}\biggl[\frac{v(X_{t})}{\rho_{t}(X_{t})}\bigg|X_{\varepsilon}=X\biggr]\biggr]$
$=$ $\,\mathbb{E}_{\mathfrak{m}}\biggl[\langle h,\,\bar{\nabla}\biggl(\rho_{\varepsilon}\,P_{t-\varepsilon}\biggl(\frac{v}{\rho_{t}}\biggr)\biggr)\rangle_{\mathcal{H}_{\mathfrak{m}}}\biggr]$
$=$ $\,\mathbb{E}_{\mathfrak{m}}\biggl[\rho_{\varepsilon}\langle h,\,\bar{\nabla}P_{t-\varepsilon}\biggl(\frac{v}{\rho_{t}}\biggr)\rangle_{\mathcal{H}_{\mathfrak{m}}}\biggr]+\mathbb{E}_{\mathfrak{m}}\biggl[\langle h,\,\bar{\nabla}\rho_{\varepsilon}\rangle_{\mathcal{H}_{\mathfrak{m}}}\,P_{t-\varepsilon}\biggl(\frac{v}{\rho_{t}}\biggr)\biggr]\,.$
Next, we can apply the Dominated Convergence Theorem since the commutation estimate Section 3 guarantees that
$\mathbb{E}_{\mathfrak{m}}\biggl[\rho_{\varepsilon}\left|\langle h,\,\bar{\nabla}P_{t-\varepsilon}\biggl(\frac{v}{\rho_{t}}\biggr)\rangle_{\mathcal{H}_{\mathfrak{m}}}\right|\biggr]=$ $\,e^{-\frac{t-\varepsilon}{2}}\mathbb{E}_{\mathfrak{m}}\biggl[\rho_{t}\left|\langle h,\,\bar{\nabla}\biggl(\frac{v}{\rho_{t}}\biggr)\rangle_{\mathcal{H}_{\mathfrak{m}}}\right|\biggr]$
$\leq$ $\,\mathbb{E}_{\overrightarrow{\mathrm{P}}_{t}}\biggl[\left|\langle h,\,\bar{\nabla}v\rangle_{\mathcal{H}_{\mathfrak{m}}}\right|\biggr]+\mathbb{E}_{\mathfrak{m}}\biggl[\frac{\left|\Gamma(\rho_{t},\,u\right|}{\rho_{t}}\biggr]$
and that
$\mathbb{E}_{\mathfrak{m}}\biggl[\left|\langle h,\,\bar{\nabla}\rho_{\varepsilon}\rangle_{\mathcal{H}_{\mathfrak{m}}}\,P_{t-\varepsilon}\biggl(\frac{v}{\rho_{t}}\biggr)\right|\biggr]=e^{\frac{t-\varepsilon}{2}}\mathbb{E}_{\mathfrak{m}}\biggl[\left|\langle h,\,\frac{\bar{\nabla}\rho_{t}}{\rho_{t}}\rangle_{\mathcal{H}_{\mathfrak{m}}}\,v\right|\biggr]\leq e^{\frac{t}{2}}\mathbb{E}_{\mathfrak{m}}\biggl[\frac{\left|\Gamma(\rho_{t},\,u)\right|}{\rho_{t}}\,|v|\biggr]\,$,
which are both uniformly bounded in $\varepsilon>0$. Therefore, from the Dominated Convergence Theorem and again from the properties of the divergence operator $\delta$ (cf. [34, Lemma 1.3.1 and Proposition 1.3.3]) we may conclude that
$\lim_{\varepsilon\downarrow 0}\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl[I(h)(X_{\varepsilon})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr]=\mathbb{E}_{\mathfrak{m}}\biggl[\langle h,\,\bar{\nabla}\biggl(\rho_{0}\,P_{t}\biggl(\frac{v}{\rho_{t}}\biggr)\biggr)\rangle_{\mathcal{H}_{\mathfrak{m}}}\biggr]=\mathbb{E}_{\overrightarrow{\mathrm{P}}}\biggl[I(h)(X_{0})\,\frac{v(X_{t})}{\rho_{t}(X_{t})}\biggr]\,$,
and hence the validity of (3.4).
Let us now focus on the last term appearing in (3.3). By relying again on the commutation estimate Section 3, the self-adjointness of the semigroup and on the properties of the divergence operator $\delta$ (cf. [34, Lemma 1.3.1 and Proposition 1.3.3]), as already done before, we have
@@
$\mathbb{E}_{\mathfrak{m}}\biggl[\frac{\Gamma(\rho_{t},\,u)}{\rho_{t}}(X)\,v(X)\biggr]=$ $\,e^{-\nicefrac{{t}}{{2}}}\,\mathbb{E}_{\mathfrak{m}}[v(X)\,\mathbb{E}_{\overrightarrow{\mathrm{P}}}[I(h)(X_{0})|X_{t}=X]]$
$-e^{-t}\mathbb{E}_{\mathfrak{m}}[v(X)\,I(h)(X)]+e^{-t}\,\mathbb{E}_{\mathfrak{m}}\biggl[\frac{\Gamma(\rho_{t},\,u)}{\rho_{t}}(X)\,v(X)\biggr],$
@@
The previous proof highlights the importance of having a commutation relation between the semigroup and the Malliavin gradient. Indeed, this is the only step in the proof where we have explicitly relied on the OU nature of the forward dynamics. This suggests that the approach presented here might be generalized to different noising (diffusion) dynamics, provided a commutation relation of the form $\bar{\nabla}P_{t}=f(t)P_{t}\bar{\nabla}$ for a general $f(t)\in(0,1)$. Particularly, this further highlights the fact that the score equals a conditional expectation of $X_{0}$ (conditioned to $\overrightarrow{X}_{t}$) provided the commutation between Malliavin gradient and semigroup, and henceforth its validity might be extended beyond the OU dynamics.
@@
Lemma 2 (Lemma 1.4.2 in [34]).
For any $u\in\mathrm{L}^{1}(\mathrm{H},\mathfrak{m})$ and any $t\geq 0$ we have
$\bar{\nabla}P_{t}\,u=e^{-\nicefrac{{t}}{{2}}}P_{t}\bar{\nabla}u\,.$
Proof.
If $u\in\mathcal{PC}$ the thesis trivially follows from the explicit expressions (2.5) combined with [34, Lemma 1.4.2]. The extension to $\mathrm{L}^{1}(\mathrm{H},\mu)$ follows from the density of $\mathcal{PC}\subseteq\mathrm{L}^{1}(\mathrm{H},\mathfrak{m})$ and the closability of $\bar{\nabla}$ [33], combined with the fact that $(P_{t})_{t\geq 0}$ is a linear contraction operator on $\mathrm{L}^{1}(\mathrm{H},\mu)$ [33, Proposition 2.8.6]. ∎
Corollary 3.
@@
Claim: $\mathrm{Q}^{J}=\overleftarrow{\mathrm{P}}^{J}$. Let us start by showing that $\mathrm{Q}^{J}$ is a solution to the (finite-dimensional) martingale problem associated to the generator $\overleftarrow{L}_{s}^{J}$ defined on any test function $u\in\mathcal{C}^{\infty}_{\mathrm{b}}(\mathbb{R}^{N})$ as
(3.7) $\overleftarrow{L}_{s}^{J}u(x)=\overrightarrow{L}^{J}u(x)+\sigma_{T-s}^{-1}\sum_{j=1}^{N}\partial_{\ell_{j}}u(x)\,\mathbb{E}_{\overrightarrow{\mathrm{P}}^{J}}[X_{0}^{\ell_{j}}-e^{-\nicefrac{{T-s}}{{2}}}X_{T-s}^{\ell_{j}}|X_{T-s}^{J}=x],$
@@
$\mathbb{E}_{\mathrm{Q}^{J}}[u(X_{t})-u(X_{0})]=\mathbb{E}_{\mathrm{Q}}[u^{\mathrm{ext}}(X_{t})-u^{\mathrm{ext}}(X_{0})]$
@@
$+\int_{0}^{t}\mathbb{E}_{\mathrm{Q}}\overrightarrow{L}u^{\mathrm{ext}}(X_{s})\mathrm{d}s$
[…]
$+\int_{0}^{t}\mathbb{E}_{\mathrm{Q}}\biggl[-\frac{1}{2}\nabla u(X^{J}_{s})\cdot X^{J}_{s}+\frac{1}{2}\sum_{j=1}^{N}C_{\ell_{j}}\partial^{2}_{jj}u(X_{s}^{J})\biggr]\mathrm{d}s$
$=\int_{0}^{t}\sigma_{T-s}^{-1}\mathbb{E}_{\mathrm{Q}}\biggl[\nabla u(X^{J}_{s})\cdot\mathbb{E}_{\overrightarrow{\mathrm{P}}^{J}}[\overrightarrow{X}^{J}_{0}-e^{\nicefrac{{T-s}}{{2}}}\overrightarrow{X}_{T-s}^{J}|\overrightarrow{X}_{T-s}^{J}=X_{s}^{J}]\biggr]\mathrm{d}s$
$+\int_{0}^{t}\mathbb{E}_{\mathrm{Q}}\overrightarrow{L}^{J}u(X^{J}_{s})\mathrm{d}s\,$,
@@
4. $\Gamma$-Score approximation and sampling scheme
@@
Therefore the core of any generative model (based on our OU noising forward dynamics) boils down to learning from noising forward dynamics the conditional expectation function $\mathbb{E}_{\overrightarrow{\mathrm{P}}}[\overrightarrow{X}_{0}|\overrightarrow{X}_{T-s}=\Phi]$. Owing to the characterization of conditional expectations as $\mathrm{L}^{2}$-projections [22, Corollary 8.17], we know that for any $\sigma(\overrightarrow{X}_{T-s})$-measurable random field $Y\in\mathrm{L}^{2}(\overrightarrow{\mathrm{P}}_{T-s})$
@@
(4.1) $\min_{\theta\in\Theta}\,\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\frac{(t_{j}-t_{j{-1}})}{T}\,\|\overrightarrow{X}^{(i)}_{0}-\mathrm{v}_{T-t_{j}}^{\theta}(\overrightarrow{X}^{(i)}_{T-t_{j}})\|_{\mathrm{H}}^{2}\,$,
@@
Input: NN samples $X^{(i)}\sim\mu_{\mathrm{data}}$ from our data distribution, time horizon $T\gg 1$ large enough, partition $0=t_{0}<t_{1}<\cdots<t_{M}=T$ (usually uniformly at random or equidistributed), parametrized vector field family $\{\mathrm{v}^{\theta}:\theta\in\Theta\}$
@@ 2   Initialize $\overrightarrow{X}_{0}^{(i)}=X^{(i)}$
[…]